OpenStudy (ujjwal):

\[\int \frac{dx}{\sinh x}\] Options are: a)cosh x +c b)log\(\frac{e^x-1}{e^x+1}\) c)\(\frac{\log e^{2x}}{1+e^x}+c\) d)-cosh x +c

OpenStudy (ujjwal):

I missed +c in option b

OpenStudy (anonymous):

Clearly, options (a) and (d) are out. Integrating csch(x) gives ln(tanh(x/2)). Does that help?

OpenStudy (foolaroundmath):

\[\sinh x = \frac{e^{x}-e^{-x}}{2}\]Take e^x = t e^x dx = dt, integral now converts to \[\int\limits \frac{2}{t^{2}-1}dt\] Can you take it off from here ? Its just simple partial fractions

OpenStudy (ujjwal):

so, its option b. Thanks @FoolAroundMath

OpenStudy (ujjwal):

And thanks @QuantumModulus